# Tackling Dimensional Collapse toward Comprehensive Universal Domain Adaptation

### v2

@@ Abstract
Universal Domain Adaptation (UniDA) addresses unsupervised domain adaptation where target classes may differ arbitrarily from source ones, except for a shared subset. An important approach, partial domain matching (PDM), aligns only shared classes but struggles in extreme cases where many source classes are absent in the target domain, underperforming the most naive baseline that trains on only source data. In this work, we identify that the failure of PDM for extreme UniDA stems from dimensional collapse (DC) in target representations. To address target DC, we propose to jointly leverage the alignment and uniformity techniques in modern self-supervised learning (SSL) on the unlabeled target data to preserve the intrinsic structure of the learned representations. Our experimental results confirm that SSL consistently advances PDM and delivers new state-of-the-art results across a broader benchmark of UniDA scenarios with different portions of shared classes, representing a crucial step toward truly comprehensive UniDA.
@@ 1 Introduction
Unsupervised Domain Adaptation (UDA) (Pan & Yang, 2009) facilitates the transfer of knowledge from a labeled source domain with a known distribution to an unlabeled target domain that may have a different distribution. Arguably the simplest UDA setting is closed-set UDA (Ganin et al., 2016; Long et al., 2018; Jiang et al., 2020), which assumes that the label sets of the source domain ($\mathcal{C}_{s}$) and the target domain ($\mathcal{C}_{t}$) are identical ($\mathcal{C}_{s}=\mathcal{C}_{t}$).
@@
While PDM approaches for UniDA appear promising, we find that they have not been comprehensively tested on a broad spectrum of UniDA scenarios. In particular, most UniDA works follow the experimental protocols established by You et al. (2019) and Fu et al. (2020), as shown in Figure 1. The protocols examined the region of relatively higher portions of target-private classes and lower portions of source-private classes. Nevertheless, they did not test UniDA approaches on high portion of source-private classes or some of the boundary cases of partial DA (X-axis) and open-set DA (Y-axis). The cases of higher source-private portions, named extreme UniDA, are especially challenging yet under-explored by the community. Such scenarios naturally arise in practice when models are pre-trained on large datasets (e.g., ImageNet) and then adapted to more specialized tasks with relatively fewer categories. In fact, our careful investigation in Figure 2 surprisingly reveals that all SOTA PDM approaches for UniDA failed in extreme UniDA, falling behind the simplest baseline of training on source data only (SO). That is, existing UniDA solutions have yet to fully address the challenges of achieving robust performance across arbitrary divisions of label sets.
@@
In contrast, we explore different modern SSL techniques that capitalize on two intuitions. First, semantically similar examples should lie closer in the representation space. Second, examples in the representation space should be spread out, which avoids degenerate solutions of mapping every example to the same representation and tackles DC directly. We adopt the framework of Wang & Isola (2020), demonstrating that the former can be realized by an alignment loss, and the latter can be achieved with a uniformity loss.
@@
    We clearly identify the DC issue behind state-of-the-art PDM methods with careful ablation study and analysis, offering novel and fundamental understanding on the (extreme) UniDA problem.
@@
    We effectively resolve the DC issue by integrating underexplored modern SSL techniques into UniDA, establishing new SOTA performance on more comprehensive benchmarks that cover all UniDA scenarios.
@@ 4 Tackling Dimensional Collapse with Self-Supervised Learning
Sun et al. (2019); Xu et al. (2019) Closed DA Pretext S+T Minimize domain gap
Bucci et al. (2019, 2021) Partial DA Pretext T Learn Target Info
Ours UniDA Cont. & Non-Cont. T Tackle DC

### v3

@@ Abstract
Universal Domain Adaptation (UniDA) addresses unsupervised domain adaptation where target classes may differ arbitrarily from source ones, except for a shared subset. A widely used approach, partial domain matching (PDM), aligns only shared classes but struggles in extreme cases where many source classes are absent in the target domain, underperforming the most naive baseline that trains on only source data. In this work, we identify that the failure of PDM for extreme UniDA stems from dimensional collapse (DC) in target representations. To address target DC, we propose to use the de-collapse techniques in self-supervised learning on the unlabeled target data to preserve the intrinsic structure of the learned representations. Our experimental results confirm that SSL consistently advances PDM and delivers new state-of-the-art results across a broader benchmark of UniDA scenarios with different portions of shared classes, representing a crucial step toward truly comprehensive UniDA. Project page: https://dc-unida.github.io/
@@ 1 Introduction
While deep learning and machine learning have achieved remarkable success across a wide range of tasks, they still struggle when there is a distribution shift between the training and testing data. Unsupervised Domain Adaptation (UDA) (Pan & Yang, 2009) addresses this challenge by transferring knowledge from a labeled source domain with a known distribution to an unlabeled target domain that may follow a different distribution. Arguably the simplest UDA setting is closed-set UDA (Ganin et al., 2016; Long et al., 2018; Jiang et al., 2020), which assumes that the label sets of the source domain ($\mathcal{C}_{s}$) and the target domain ($\mathcal{C}_{t}$) are identical ($\mathcal{C}_{s}=\mathcal{C}_{t}$).
@@
While PDM methods for UniDA appear promising, they have not been comprehensively tested on a broad spectrum of UniDA scenarios. In particular, most UniDA works follow the experimental protocols established by You et al. (2019) and Fu et al. (2020), as shown in Figure 1. The protocols examined the region of relatively higher portions of target-private classes and lower portions of source-private classes. Nevertheless, they did not test UniDA methods on high portion of source-private classes or some of the boundary cases of partial DA (X-axis) and open-set DA (Y-axis). The cases of higher source-private portions, named extreme UniDA, are especially challenging yet under-explored by the community. Such scenarios naturally arise in practice when models are pre-trained on large datasets and then adapted to more specialized tasks with fewer categories. In fact, our careful investigation in Figure 2 surprisingly reveals that all SOTA PDM methods for UniDA failed in extreme UniDA, falling behind the simplest baseline of training on source data only (SO). That is, existing UniDA solutions have yet to fully address the challenges of achieving robust performance across arbitrary divisions of label sets.
@@
In contrast, we explore different SSL techniques that capitalize on two intuitions. First, semantically similar examples should lie closer in the representation space. Second, examples in the representation space should be spread out, which avoids degenerate solutions of mapping every example to the same representation and tackles DC directly. We adopt the framework of Wang & Isola (2020), demonstrating that the former can be realized by an alignment loss, and the latter can be achieved with a uniformity loss.
@@
    We identify the DC issue behind state-of-the-art PDM methods with careful ablation study and analysis, offering novel and fundamental understanding on the (extreme) UniDA problem.
@@
    We effectively resolve the DC issue by integrating underexplored SSL techniques into UniDA, establishing new SOTA performance on more comprehensive benchmarks that cover all UniDA scenarios.
@@ 4 Tackling Dimensional Collapse with Self-Supervised Learning
Sun et al. (2019); Xu et al. (2019) Closed DA Pretext S+T Reduce negative transfer
Bucci et al. (2019, 2021) Partial DA Pretext T Reduce negative transfer
Ours UniDA Cont. & Non-Cont. T De-collapse
@@ 6 Conclusion
Acknowledgments
We thank the anonymous reviewers and members of CLLab for their constructive feedback. This work is supported by the National Science and Technology Council in Taiwan via NSTC 113-2628-E-002-003 and NSTC 113-2634-F-002-008. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.